# Aligning Large Language Models with Representation Editing: A Control Perspective

### v2

@@ Abstract
Aligning large language models (LLMs) with human objectives is crucial for real-world applications. However, fine-tuning LLMs for alignment often suffers from unstable training and requires substantial computing resources. Test-time alignment techniques, such as prompting and guided decoding, do not modify the underlying model, and their performance remains dependent on the original model’s capabilities. To address these challenges, we propose aligning LLMs through representation editing. The core of our method is to view a pre-trained autoregressive LLM as a discrete-time stochastic dynamical system. To achieve alignment for specific objectives, we introduce external control signals into the state space of this language dynamical system. We train a value function directly on the hidden states according to the Bellman equation, enabling gradient-based optimization to obtain the optimal control signals at test time. Our experiments demonstrate that our method outperforms existing test-time alignment techniques while requiring significantly fewer resources compared to fine-tuning methods.
@@ 2.1 Large Language Model Alignment
Alignment through Fine-tuning:
RLHF has been a popular method in LLM alignment stiennon2020learning ; zhu2023principled ; touvron2023llama . While effective, RLHF entails a complex process that involves training multiple models and continuously sampling from the LM policy during the learning loop. DPO rafailov2023direct simplifies the RLHF framework by using a direct optimization objective derived from Proximal Policy Optimization (PPO; schulman2017proximal ), reducing the process to supervised training of the policy model alone. However, DPO is memory-intensive and resource-demanded as it requires managing two policies simultaneously. Contrastive Preference Optimization (CPO; xu2024contrastive ) mitigates these challenges by utilizing a uniform reference model, which not only reduces memory requirements but also enhances training efficiency. Alternative methods such as yuan2023rrhf ; song2023preference simplify model management and parameters tuning in the RLHF framework by adopting a supervised fine-tuning (SFT) approach. Additionally, RSO liu2023statistical and RAFT dong2023raft employ rejection sampling to refine the alignment process. RSO focuses on estimating the optimal policy more accurately, while RAFT uses high-quality samples for iterative fine-tuning of the policy model.
@@
Test time alignment:
@@ 4.3 Training of Value Function
Parameterization of the Value Function
The simplest approach is to add control signals only to the logit $o_{t}$. In this case, we can directly use a simple neural network as the value function. If we want to incorporate the attention key-value pairs $h_{t}$ , we need to handle the varying size of the input. To achieve this, we can initialize a vector and compute an attention weight by taking the dot product with the keys to aggregate all value embeddings. Then, we concatenate the aggregated value embedding with $o_{t}$ and input it into a neural network. In practice, we find that parameterizing the value function as a two- or three-layer neural network is sufficient for achieving good empirical performance.
@@ 5.1 Experimental Setup
We evaluate our method on the HH-RLHF bai2022training dataset, which is the most widely used dataset for LLM alignment. This dataset is used to improve the AI assistant’s helpfulness and harmlessness, comprising 161,000 training samples and 8,550 test samples. Each sample contains a prompt and two responses with one preferred over another. For the base model, we adopt Vicuna-7B chiang2023vicuna and Falcon-7B-Instruct almazrouei2023falcon as the instructed fine-tuned AI assistant. We evaluate these models by generating text responses based on test prompts from of HH-RLHF. For the reward model, we use a publicly available one that employs LLaMA-7B [^1] https://huggingface.co/argsearch/llama-7b-rm-float32 as the backbone, trained on HH-RLHF using the pairwise reward loss ouyang2022training . We train the value network on the last layer of the hidden states $o_{t}$, and at test time, we add control signals only to this layer. For future studies, we can also explore adding controls to the attention key-value pairs $h_{t}$ which should further improve the performance.
Following khanov2024args , we leverage Diversity, Coherence, Average Reward, and Win Rate as our evaluation metrics. Diversity measures the frequency of repeated n-grams in generated text. The diversity score for a given response $\mathbf{y}$ is represented as $\prod_{n=2}^{4}\frac{\text{unique n-grams}(\mathbf{y)}}{\text{total n-grams}(\mathbf{y})}$. A higher diversity score suggests a broader vocabulary range in text generation. Coherence calculates the cosine similarity between the embeddings of the prompt and its continuation. We use the pre-trained SimCSE sentence embedding model, following the approach outlined in yixuansu2022 , to obtain these embeddings. Average Reward is the mean of the rewards evaluated by the reward model across all responses corresponding to the test prompts. Win Rate is the rate at which the model’s response is rated better than the preferred response in the dataset. Following khanov2024args ; chiang2023vicuna , we use GPT-4 as the judge, having it review and score two responses to the same prompt on a scale from 1 to 10. We provide explicit instructions to assess the responses based on criteria such as helpfulness, harmlessness, relevance, accuracy, and insightfulness. The detailed prompt is provided in Appendix D. We randomly sample 300 prompts from the test set of HH-RLHF for the GPT-4 evaluation. To mitigate position bias, we randomize the order in which we present the generated responses to GPT-4, as in zheng2023judging .
@@
5.2 Baselines
Prompt Engineering: In this method, we instruct the model within the prompt to provide responses that are more helpful and harmless touvron2023llama . Controlled Decoding (CD): During the decoding process of LLMs, this method combines token probabilities with reward scores. We consider two versions. The first version khanov2024args directly uses a reward model trained on human preference data, requiring the tokenization strategies of both the reward model and the base model to be the same. The second version mudgal2023controlled trains a prefix scorer to predict the expected reward from partially generated responses. We refer to it as CD prefix. Static Representation Editing (RE): Following li2024inference , we first train a linear regression layer on the hidden state of an LLM, after feeding it the prompt, to predict the expected reward. At test time, we shift the activation space along the direction of the weights of the linear layer. Unlike our method, this approach adds a fixed vector to the representation space along the generation trajectory.
Backbone Model Diversity $\uparrow$ Coherence $\uparrow$ Average Reward $\uparrow$ Win Rate (%) $\uparrow$ Inference time (hour)
Vicuna 7B Base 0.816 0.568 5.894 57.6 0.60
Prompting 0.817 0.570 5.913 66.0 0.69
@@ 5.2 Baselines
Table 1: Performance comparison between Re-Control and other test-time alignment approaches on HH-RLHF. The win rate is evaluated by GPT-4 as the rate at which the model’s response is rated better than the preferred response in the dataset. Note that CD khanov2024args requires the base model to have the same tokenization strategy as the reward model.
@@ 5.3 Experimental Results
Table 1 shows the performance of all the methods. Our findings can summarized as follows: (1) Re-Control achieves the highest alignment score in terms of the win rate evaluated by GPT-4. Moreover, it maintains generation quality, as measured by diversity and coherence. While controlled decoding achieves the best average reward on Falcon-7B, Re-Control outperforms it in terms of the other three metrics. This suggests that controlled decoding may encounter reward overoptimization. (2) The strongest baseline is controlled decoding. However, controlled decoding is 20 times slower than Re-Control. This is because controlled decoding needs to evaluate multiple candidate tokens and perform forward passes through the entire reward model repeatedly, while Re-Control only requires optimization through a value function that is a two- or three-layer neural network, making it much faster. (3) Combining prompt engineering with Re-Control can further improve alignment performance in terms of both average reward and GPT-4 evaluation. Specifically, it outperforms the strongest baseline by 5.5% and 9.7% in terms of the GPT-4 win rate. In contrast, controlled decoding with prompting shows only marginal improvements. This might be because Re-Control perturbs the activation space of the LLM, which is more flexible than merely changing the final token probability. (4) Re-Control significantly outperforms static representation editing by 11.7% on Vicuna-7B and 9.7% on Falcon-7B. This is because Re-Control dynamically adjusts the representation during the autoregressive generation process, offering more control. In contrast, static representation editing applies a fixed shift, which is more rigid.
@@ 6.1 Comparison with Training-time Alignment
Figure 3: Comparison with fine-tuning methods using Vicuna-7B as the base model.
In the previous section, we compared Re-Control with test-time alignment methods that do not require extensive computing resources. This feature is crucial when we need the model to quickly adapt to different requirements, as it only involves training a simple value network with just two or three layers. In this subsection, we further compare Re-Control with fine-tuning based approaches. Figure 3 shows the comparison between Re-Control, Proximal Policy Optimization (PPO), and Direct Preference Optimization (DPO) rafailov2023direct . All the models use Vicuna-7B as the base model. The training details for PPO and DPO are provided in C. We observe that Re-Control achieves a higher GPT-4 win rate and average reward compared to both PPO and DPO. Furthermore, Re-Control also outperforms these methods in terms of diversity and coherence. Overall, the results indicate that our approach is a competitive alternative to fine-tuning methods.
@@ 6.2 Generalization to a new input distribution
An important question is how our method can generalize to a new input distribution different from the value function is trained on. To investigate this question, we further test on a out-of-distribution (OOD) dataset HarmfulQA bhardwaj2023red . The test split of HarmfulQA contains harmful questions to evaluate language model performance against red-teaming attempts. We focus on the GPT-4 evaluation since the reward model will not be accurate for the OOD data. We compare Re-Control + promoting with other test-time alignment methods + prompting. Figure 4 presents the results. As illustrated, Re-Control + Prompting achieves the highest performance in terms of the GPT-4 win rate on both Vicuna-7B and Falcon-7B. This is an important ability especially when we want to deploy the LLM in the open world.
@@ Appendix A Limitations and Future Work
We discuss limitations and possible extensions of Re-Control. (1) Injecting inductive bias into the control policy. In our current work, we only train a value function on the last layer of the model’s hidden space. However, we can follow the approach in li2024inference , first training multiple value functions on all intermediate hidden layers and then selecting the layer that achieves the best accuracy on the validation set. Additionally, we can draw from the methods in geiger2024finding ; wu2024reft ; wei2024assessing to perturb only a low-rank subspace of the representation space. (2) Multi-objective aligment. In the current paper, we consider the objective from a single reward model. However, in practice, alignment may involve multiple, potentially conflicting objectives. It would be interesting to leverage multi-objective optimization techniques gunantara2018review at test time to obtain a Pareto frontier in the representation space for such settings. (3) More advanced training algorithm. Currently, we train the value function using a simple one-iteration policy iteration method. It would be interesting to explore whether increasing the number of iterations could further improve the training of the value function. Additionally, we can consider using algorithms for training the value function that provide provable convergence wang2022convergent .
@@ C.2 HH-RLHF
We evaluate our method on the HH-RLHF bai2022training dataset, which is the most widely used dataset for LLM alignment. This dataset is used to improve the AI assistant’s helpfulness and harmlessness, comprising 161,000 training samples and 8,550 test samples. Each sample contains a prompt and two responses with one preferred over another. For the base model, we adopt Vicuna-7B [^2] https://huggingface.co/lmsys/vicuna-7b-v1.5 chiang2023vicuna and Falcon-7B-Instruct [^3] https://huggingface.co/tiiuae/falcon-7b almazrouei2023falcon as the instructed fine-tuned AI assistant. We evaluate these models by generating text responses based on test prompts from of HH-RLHF. Following the standard practice, we limit the maximum lengths of the prompt and generated continuation to $2,048$ and $128$ tokens, respectively.
For the reward model, we use a publicly available one that employs LLaMA-7B [^4] https://huggingface.co/argsearch/llama-7b-rm-float32 as the backbone, trained on HH-RLHF using the pairwise reward loss ouyang2022training .
@@
To train the value function of Re-Control, we adopt the Adam optimizer kingma2014adam . The training hyperparameters of the value networks are summarized in Table  3.
@@
Vicuna Number of epochs 100
@@
Falcon Number of epochs 100
@@
Table 3: Summary of the hyperparameters used in training the value function of Re-Control.
@@
Vicuna Step size 0.5
@@
Maximum lengths of genearted continuation 128
Falcon Step size 0.2
@@
Maximum lengths of genearted continuation 128
Table 4: Summary of hyperparameters of Re-Control at test time.
@@
We first train a linear regression layer on the hidden state of a large language model (LLM) after feeding the prompt, to predict the expected reward as in li2024inference . For a fair comparison, we use the same hidden state layer as Re-Control. At test time, we shift the activation space along the direction of the weights using an intervention strength parameter $\alpha$, which is selected based on the validation set. The hyperparameters used during the training and testing stages are summarized in Table 5.
@@
We use the codebase [^5] https://github.com/deeplearning-wisc/args from khanov2024args . We employ the default hyperparameters suggested in the paper and repository. The number of candidates to rank with the reward model is set to 10, and the weight controlling the tradeoff between the LLM text objective and the reward is 1. For controlled decoding with the value function, we stack the value function of Re-Control on top of the hidden state of the LLM as the prefix scorer, ensuring a fair comparison with our method.
@@
Vicuna Number of epochs 100
@@
Falcon Number of epochs 100
@@
Table 5: Summary of hyperparameters of static representation editing
@@
Vicuna Max number of PPO update steps 10000
@@
Vicuna Max number of training steps 10000
@@
C.3 HarmfulQA
This dataset [^6] https://huggingface.co/datasets/declare-lab/HarmfulQA contains 1,960 harmful questions specifically designed to evaluate the performance of language models. Additionally, it includes a conversation set comprising 9,536 harmless conversations and 7,356 harmful conversations for model alignment purposes. For our experiments, we focused solely on the evaluation portion of the HarmfulQA dataset to test the performance of our method.

### v3

@@ Abstract
Aligning large language models (LLMs) with human objectives is crucial for real-world applications. However, fine-tuning LLMs for alignment often suffers from unstable training and requires substantial computing resources. Test-time alignment techniques, such as prompting and guided decoding, do not modify the underlying model, and their performance remains dependent on the original model’s capabilities. To address these challenges, we propose aligning LLMs through representation editing. The core of our method is to view a pre-trained autoregressive LLM as a discrete-time stochastic dynamical system. To achieve alignment for specific objectives, we introduce external control signals into the state space of this language dynamical system. We train a value function directly on the hidden states according to the Bellman equation, enabling gradient-based optimization to obtain the optimal control signals at test time. Our experiments demonstrate that our method outperforms existing test-time alignment techniques while requiring significantly fewer resources compared to fine-tuning methods. Our code is available at https://github.com/Lingkai-Kong/RE-Control.
@@ 2.1 Large Language Model Alignment
Alignment through Fine-tuning.
RLHF has been a popular method in LLM alignment stiennon2020learning ; zhu2023principled ; touvron2023llama . While effective, RLHF entails a complex process that involves training multiple models and continuously sampling from the LM policy during the learning loop. DPO rafailov2023direct simplifies the RLHF framework by using a direct optimization objective derived from Proximal Policy Optimization (PPO; schulman2017proximal ), reducing the process to supervised training of the policy model alone. However, DPO is memory-intensive and resource-demanded as it requires managing two policies simultaneously. Contrastive Preference Optimization (CPO; xu2024contrastive ) mitigates these challenges by utilizing a uniform reference model, which not only reduces memory requirements but also enhances training efficiency. Alternative methods such as yuan2023rrhf ; song2023preference simplify model management and parameter tuning in the RLHF framework by adopting a supervised fine-tuning (SFT) approach. Additionally, RSO liu2023statistical and RAFT dong2023raft employ rejection sampling to refine the alignment process. RSO focuses on estimating the optimal policy more accurately, while RAFT uses high-quality samples for iterative fine-tuning of the policy model.
@@
Test time alignment.
@@ 4.4 Test-time Intervention
Parameterization of the Value Function. Rigorously, policy iteration requires the input to the value function to be the full state, but it does not require the control signals to be added to the full state. This means we can train the value function on the full state and backpropagate through it with respect to partial inputs at test time. The simplest approach is to add control signals only to the logit $o_{t}$. In this case, we find that training a two- or three-layer neural network using $o_{t}$ as the input is already sufficient for achieving good empirical performance. If we want to further incorporate the attention key-value pairs $h_{t}$ in the input, we need to address the input’s varying size. To achieve this, we can initialize a vector and compute an attention weight by taking the dot product with the keys to aggregate all value embeddings. Then, we concatenate the aggregated value embedding with $o_{t}$ and input it into a neural network.
@@ 5.1 Experimental Setup
We evaluate our method on the HH-RLHF bai2022training and Stanford SHP (SHPpmlr-v162-ethayarajh22a datasets, which are popular for LLM alignment. These two datasets are used to improve the AI assistant’s helpfulness and harmlessness. Each sample in the datasets contains a prompt and two responses with one preferred over another. For the base model, we adopt Vicuna-7B chiang2023vicuna , Falcon-7B almazrouei2023falcon and Llama3-8B dubey2024llama as the instructed fine-tuned AI assistant. We evaluate these models by generating text responses based on test prompts of HH-RLHF and SHP. For reproducibility, we use publicly available reward models [^1] HH-RLHF: https://huggingface.co/argsearch/llama-7b-rm-float32  [^2] SHP: https://huggingface.co/openbmb/UltraRM-13b. We train the value network on the last layer of the hidden states $o_{t}$, and at test time, we add control signals only to this layer. For future studies, we can also explore adding controls to the attention key-value pairs $h_{t}$ which should further improve the performance.
Following khanov2024alignment , we leverage Diversity, Coherence, Average Reward, and Win Rate as our evaluation metrics. Diversity measures the frequency of repeated n-grams in generated text. The diversity score for a given response $\mathbf{y}$ is represented as $\prod_{n=2}^{4}\frac{\text{unique n-grams}(\mathbf{y)}}{\text{total n-grams}(\mathbf{y})}$. A higher diversity score suggests a broader vocabulary range in text generation. Coherence calculates the cosine similarity between the embeddings of the prompt and its continuation. We use the pre-trained SimCSE sentence embedding model, following the approach outlined in yixuansu2022 , to obtain these embeddings. Average Reward is the mean of the rewards evaluated by the reward model across all responses corresponding to the test prompts. Win Rate is the rate at which the model’s response is rated better than the preferred response in the dataset. Following khanov2024alignment ; chiang2023vicuna , we use GPT-4 as the judge, having it review and score two responses to the same prompt on a scale from 1 to 10. We provide explicit instructions to assess the responses based on criteria such as helpfulness, harmlessness, relevance, accuracy, and insightfulness. The detailed prompt is provided in Appendix D. We randomly sample 300 prompts from the test set of HH-RLHF for the GPT-4 evaluation. To mitigate position bias, we randomize the order in which we present the generated responses to GPT-4, as in zheng2023judging . Additionally, we also present the Inference Time in hour under batch size of 32.
@@
Dataset Backbone Model Diversity $\uparrow$ Coherence $\uparrow$ Average Reward $\uparrow$ Win Rate (%) $\uparrow$ Inference time (hour)
HH-RLHF Vicuna-7B Base 0.816 0.568 5.894 57.6 0.60
@@
SHP Vicuna-7B Base 0.845 0.657 -5.68 40.3 0.13
Static RE 0.848 0.652 -5.65 49.3 0.15
CD 0.845 0.655 -5.65 55.6 22.16
CD prefix 0.838 0.660 -5.62 41.0 14.15
Ours 0.849 0.652 -5.38 58.0 0.21
Prompting 0.847 0.570 -4.83 56.6 0.13
CD prefix + Prompting 0.842 0.574 -4.88 56.3 14.32
Ours + Prompting 0.854 0.571 -4.63 63.6 0.23
Llama3-8B Base 0.878 0.672 -4.64 56.3 0.13
Static RE 0.875 0.674 -4.49 57.0 0.15
CD N/A N/A N/A N/A N/A
CD prefix 0.862 0.685 -4.41 64.0 12.16
Ours 0.883 0.669 -4.39 71.0 0.21
Prompting 0.891 0.605 -4.45 59.6 0.13
CD prefix + Prompting 0.872 0.603 -4.25 68.0 12.74
Ours + Prompting 0.893 0.605 -4.14 77.0 0.24
Table 1: Performance comparison between Re-Control and other test-time alignment approaches. The win rate is evaluated by GPT-4 as the rate at which the model’s response is rated better than the preferred response in the dataset. Note that CD khanov2024alignment requires the base model to have the same tokenization strategy as the reward model.
5.2 Baselines
Prompt Engineering: In this method, we instruct the model within the prompt to provide responses that are more helpful and harmless touvron2023llama . Controlled Decoding (CD): During the decoding process of LLMs, this method combines token probabilities with reward scores. We consider two versions. The first version khanov2024alignment directly uses a reward model trained on human preference data, requiring the tokenization strategies of both the reward model and the base model to be the same. The second version mudgal2023controlled trains a prefix scorer to predict the expected reward from partially generated responses. We refer to it as CD prefix. Static Representation Editing (RE): Following li2024inference , we first train a linear regression layer on the hidden state of an LLM, after feeding it the prompt, to predict the expected reward. At test time, we shift the activation space along the direction of the weights of the linear layer. Unlike our method, this approach adds a fixed vector to the representation space along the generation trajectory.
@@ 5.3 Experimental Results
Table 1 shows the performance of all the methods. Our findings can summarized as follows: (1) Re-Control achieves the highest alignment score in terms of the win rate evaluated by GPT-4. Moreover, it maintains generation quality, as measured by diversity and coherence. (2) The strongest baseline is controlled decoding. However, controlled decoding is 20 times slower than Re-Control. This is because controlled decoding needs to evaluate multiple candidate tokens and perform forward passes through the entire reward model repeatedly, while Re-Control only requires optimization through a value function that is a two- or three-layer neural network, making it much faster. Additionally, CD khanov2024alignment lacks support for batch generation, making it even slower. (3) Combining prompt engineering with Re-Control can further improve alignment performance in terms of both average reward and GPT-4 evaluation. Specifically, it outperforms the strongest baseline by {7.6%, 19.0%, 12.4%, 13.2%} in terms of the GPT-4 win rate. In contrast, controlled decoding with prompting shows only marginal improvements. This might be because Re-Control perturbs the activation space of the LLM, which is more flexible than merely changing the final token probability. (4) Re-Control significantly outperforms static representation editing by {17.6%, 10.2% 17.6%, 24.6%}. This is because Re-Control dynamically adjusts the representation during the autoregressive generation process, offering more control. In contrast, static representation editing applies a fixed shift, which is more rigid.
@@ 6.1 Comparison with Training-time Alignment
In the previous section, we compared Re-Control with test-time alignment methods that do not require extensive computing resources. This feature is crucial when we need the model to quickly adapt to different requirements, as it only involves training a simple value network with just two or three layers. In this subsection, we further compare Re-Control with fine-tuning based approaches with LoRA hu2021lora . Figure 3 shows the comparison between Re-Control, Proximal Policy Optimization (PPO), and Direct Preference Optimization (DPO) rafailov2023direct .
Figure 3: Comparison with LoRa-based fine-tuning methods using Vicuna-7B as the base model on HH-RLHF.
All the models use Vicuna-7B as the base model and we test them on HH-RLHF. The training details for LoRa-based PPO and DPO are provided in C. Overall, the results indicate that our approach is a competitive alternative to LoRa-based fine-tuning methods. Similar findings have also been reported in the controlled decoding literature khanov2024alignment . Overall, for users prioritizing real-time inference, amortizing computation during the training process remains preferable. However, for those without resources for fine-tuning, test-time alignment is a more practical choice, as it easily adapts to different alignment objectives, albeit with increased inference time.
@@ 6.2 Generalization to a new input distribution
An important question is how our method can generalize to a new input distribution different from the value function is trained on. To investigate this question, we further test on a out-of-distribution (OOD) dataset HarmfulQA bhardwaj2023red with the value function trained on HH-RLHF. The test split of HarmfulQA contains harmful questions to evaluate language model performance against red-teaming attempts. We focus on the GPT-4 evaluation since the reward model will not be accurate for the OOD data. We compare Re-Control + Prompting with other test-time alignment methods + Prompting. Figure 4 presents the results. As illustrated, Re-Control + Prompting achieves the highest performance in terms of the GPT-4 win rate on both Vicuna-7B and Falcon-7B. This is an important ability especially when we want to deploy the LLM in the open world.
@@ 6.3 Hyperparameter Study
6.4 Inference Time Analysis
Figure 6: Inference time comparison under different batch sizes. Note that CD does not support batch generation.
Figure 7: Compute-performance tradeoff at test time on HH-RLHF using Vicuna-7B as the backbone. Note that CD does not support batch generation.
We provide additional analysis of the inference time. Figure 7 presents the inference time across different batch sizes. As shown, increasing the batch size reduces the discrepancy between Re-Control and the base model, becoming negligible at a batch size of 32. Figure 7 illustrates the compute-performance tradeoff between Re-Control and CD. For Re-Control, we vary the number of iterations when optimizing through the value function at test time, while for CD, we adjust the number of candidate tokens. As shown, the performance of RE-Control initially increases with more computing time but eventually decreases. This decline occurs because a large number of iterations at test time can lead to reward hacking, reducing the quality of the generated sentences. As discussed in Section 6.3, this hyperparameter can be selected based on the validation set. Since CD does not support batch generation, its inference speed is significantly slower than Re-Control. Even when RE-Control does not use batch generation, it outperforms CD when using the same computing resources. For example, when the inference time is around $155$ minutes, the win rate of Re-Control is $75\%$, while CD is only $62\%$.
@@ 7 Conclusion, Limitations and Future Work
Acknowledgments and Disclosure of Funding
We thank the anonymous reviewers for their helpful comments. This work was supported in part by NSF IIS-2008334, IIS-2106961, IIS-2403240, and CAREER IIS-2144338, Schmidt Sciences AI2050 Fellowship and computing resources from Georgia Tech.
@@ Appendix A Limitations and Future Work
We discuss limitations and possible extensions of Re-Control. (1) Injecting inductive bias into the control policy. In our current work, we only train a value function on the last layer of the model’s hidden space. However, we can follow the approach in [35], first training multiple value functions on all intermediate hidden layers and then selecting the layer that achieves the best accuracy on the validation set. Additionally, we can draw from the methods in [23, 68, 64] to perturb only a low-rank subspace of the representation space. (2) Multi-objective aligment. In the current paper, we consider the objective from a single reward model. However, in practice, alignment may involve multiple, potentially conflicting objectives. It would be interesting to leverage multi-objective optimization techniques [24] at test time to obtain a Pareto frontier in the representation space for such settings. (3) More advanced training algorithm. Currently, we train the value function using a simple one-iteration policy iteration method. It would be interesting to explore whether increasing the number of iterations could further improve the training of the value function. Additionally, we can consider using algorithms for training the value function that provide provable convergence [63].
@@ C.2 HH-RLHF
We evaluate our method on the HH-RLHF [5] dataset, which is the most widely used dataset for LLM alignment. This dataset is used to improve the AI assistant’s helpfulness and harmlessness, comprising 161,000 training samples and 8,550 test samples. Each sample contains a prompt and two responses with one preferred over another. For the base model, we adopt Vicuna-7B [^3] https://huggingface.co/lmsys/vicuna-7b-v1.5 [12] and Falcon-7B-Instruct [^4] https://huggingface.co/tiiuae/falcon-7b [3] as the instructed fine-tuned AI assistant. We evaluate these models by generating text responses based on test prompts from of HH-RLHF. Following the standard practice, we limit the maximum lengths of the prompt and generated continuation to $2,048$ and $128$ tokens, respectively.
For the reward model, we use a publicly available one that employs LLaMA-7B [^5] https://huggingface.co/argsearch/llama-7b-rm-float32 as the backbone, trained on HH-RLHF using the pairwise reward loss [45].
@@
To train the value function of Re-Control, we adopt the Adam optimizer [32]. The training hyperparameters of the value networks are summarized in Table  3.
@@
Vicuna-7B Number of epochs 100
@@
Falcon-7B Number of epochs 100
@@
Table 3: Summary of the hyperparameters used in training the value function of Re-Control on HH-RLHF.
@@
Vicuna-7B Step size 0.5
@@
Maximum lengths of generated continuation 128
Falcon-7B Step size 0.2
@@
Maximum lengths of generated continuation 128
Table 4: Summary of hyperparameters of Re-Control at test time on HH-RLHF.
@@
We first train a linear regression layer on the hidden state of a large language model (LLM) after feeding the prompt, to predict the expected reward as in [35]. For a fair comparison, we use the same hidden state layer as Re-Control. At test time, we shift the activation space along the direction of the weights using an intervention strength parameter $\alpha$, which is selected based on the validation set. The hyperparameters used during the training and testing stages are summarized in Table 5.
@@
We use the codebase [^6] https://github.com/deeplearning-wisc/args from [31]. We employ the default hyperparameters suggested in the paper and repository. The number of candidates to rank with the reward model is set to 10, and the weight controlling the tradeoff between the LLM text objective and the reward is 1. For controlled decoding with the value function, we stack the value function of Re-Control on top of the hidden state of the LLM as the prefix scorer, ensuring a fair comparison with our method.
@@
Vicuna-7B Number of epochs 100
@@
Falcon-7B Number of epochs 100
@@
Table 5: Summary of hyperparameters of static representation editing on HH-RLHF
@@
Vicuna-7B Max number of PPO update steps 10000
@@
Vicuna-7B Max number of training steps 10000
@@
C.3 Standford SHP
To further evaluate our method, we utilized the Standford SHP (SHP[20] dataset. This dataset comprises 385,000 collective human preferences across diverse subject areas, ranging from cooking to legal advice. Each data point consists of a Reddit post with a question or instruction, and two top-level comments, one of which has been rated as more helpful by Reddit users. The dataset is divided into 349,000 training samples, 18,400 validation samples, and 18,400 test samples, enabling robust evaluation of our approach. For the base model, we adopt Vicuna-7B [^7] https://huggingface.co/lmsys/vicuna-7b-v1.5 [12] and Llama3-8B-Instruct [^8] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct [2] as the instructed fine-tuned AI assistant. We evaluate these models by generating text responses based on 1,000 random sampled test prompts from of SHP. Following the standard practice, we limit the maximum lengths of the prompt and generated continuation to $2,048$ and $128$ tokens, respectively.
For the reward model, we use UltraRM-13B [^9] https://huggingface.co/openbmb/UltraRM-13b [^10] During the rebuttal stage, we used another publicly available reward model, but further investigation revealed it to be unreliable, so we opted to use a different reward model., which is trained on Anthropic HH-RLHF, Standford SHP, and Summarization.
Re-Control.
When constructing the training dataset for the value function, we sample only one response for each training prompt of Standford SHP, i.e., $M=1$. For both Vicuna-7B and Llama3-8B-Instruct, we train the value network on the last layer of the hidden states $o_{t}$, and at test time, we add control signals only to this layer. For future studies, we can also explore adding controls to the attention key-value pairs $h_{t}$ which should further improve the performance.
For both Vicuna-7B and Llama3-8B-Instruct, the value function is a two-layer network with a hidden dimension of 4096.
To train the value function of Re-Control, we adopt the Adam optimizer [32]. The training hyperparameters of the value networks are summarized in Table  8.
Backbone Parameters Value
Vicuna-7B Number of epochs 100
Learning rate $1*10^{-4}$
batch size 512
Floating point format fp16 (Half-precision)
Number of Layers 2
Hidden Dimension 4096
Llama3-8B Number of epochs 100
Learning rate $1*10^{-4}$
batch size 512
Floating point format fp16 (Half-precision)
Number of Layers 2
Hidden Dimension 4096
Table 8: Summary of the hyperparameters used in training the value function of Re-Control on SHP.
The step size $\alpha$ and number of updates $n$ are selected on the validation set to maximize the sum of coherence, diversity, and average reward. The inference parameters are summarized in Table 9.
Backbone Parameters Value
Vicuna-7B Step size 1.0
Number of updates 50
batch size 32
Floating point format fp16 (Half-precision)
Maximum lengths of the prompt 2048
Maximum lengths of generated continuation 128
Llama3-8B Step size 1.0
Number of updates 30
batch size 32
Floating point format fp16 (Half-precision)
Maximum lengths of the prompt 2048
Maximum lengths of generated continuation 128
Table 9: Summary of hyperparameters of Re-Control at test time on SHP.
Prompting engineering.
"A question from a curious user and an answer from an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions." + Original prompt
Static representation editing.
We first train a linear regression layer on the hidden state of a large language model (LLM) after feeding the prompt, to predict the expected reward as in [35]. For a fair comparison, we use the same hidden state layer as Re-Control. At test time, we shift the activation space along the direction of the weights using an intervention strength parameter $\alpha$, which is selected based on the validation set. The hyperparameters used during the training and testing stages are summarized in Table 10.
Controlled Decoding.
We use the codebase [^11] https://github.com/deeplearning-wisc/args from [31]. We employ the default hyperparameters suggested in the paper and repository. The number of candidates to rank with the reward model is set to 10, and the weight controlling the tradeoff between the LLM text objective and the reward is 1. For controlled decoding with the value function, we stack the value function of Re-Control on top of the hidden state of the LLM as the prefix scorer, ensuring a fair comparison with our method.
Backbone Parameters Value
Vicuna-7B Number of epochs 100
Learning rate $1*10^{-4}$
Training batch size 512
Testing batch size 32
Intervention strength 2.5
Llama3-8B Number of epochs 100
Learning rate $1*10^{-3}$
Training batch size 512
Testing batch size 32
Intervention strength 1.5
Table 10: Summary of hyperparameters of static representation editing on SHP.
C.4 HarmfulQA
This dataset [^12] https://huggingface.co/datasets/declare-lab/HarmfulQA contains 1,960 harmful questions specifically designed to evaluate the performance of language models. Additionally, it includes a conversation set comprising 9,536 harmless conversations and 7,356 harmful conversations for model alignment purposes. For our experiments, we focused solely on the evaluation portion of the HarmfulQA dataset to test the performance of our method. We use the value function trained on HH-RLHF to evaluate the out-of-distribution performance.